# Visual Thoughts: A Unified Perspective of Understanding Multimodal Chain-of-Thought

### v1

@@ Appendix A Experiment Setup
We conduct all our experiments using four models, including LLaVA-1.5 [21], Qwen2-VL [42], GPT-4o-mini [32], and GPT-4o [32]. For GPT models, we adjust the temperature parameter within [0,2]; for the open-source models, we adjust the temperature parameter within the range [0,2]. In addition, all open source models complete inference on 2 A6000 48G.
@@ B.2 Image Classification
In this section we investigate how the complexity of an image’s textual description influences the effectiveness of visual reasoning. We first prompt the model to produce captions that are as precise as possible, and we quantify description difficulty by the number of tokens in each caption. Consistent with commonsense expectations and our manual annotations, images depicting spot-the-difference scenarios are the most challenging to describe and are therefore labeled Hard-to-Describe. Tangram-related images present a moderate level of difficulty and are labeled Medium-to-Describe. By contrast, images showing mathematical functions detected through optical character recognition (OCR) require the fewest tokens and are labeled Easy-to-Describe.
B.3 More Results
Beyond the exploration in Section 3, we further investigate the effectiveness of incorporating visual thought into pure-text problems that require spatial imagination. In this setting, visual thought serves to convey visual information from the imagined spatial domain that is necessary for solving the textual problems. Specifically, we select tasks including chess, function, and graph from IsoBench and use Qwen2-VL as the experimental model. We employ VisualSketchpad[15] to generate reasoning processes that include image-form visual thought, and generate the caption for Caption-Only based on the input textual mathematical expression, keeping all other settings the same.
@@ Appendix C The Categories of Visual Thoughts
In this part of the experiment, we assess the performance of four types of visual thoughts using four models, including LLaVA-1.5 [21], Qwen2-VL [42], GPT-4o-mini [32], and GPT-4o [32].
@@ C.1 Prompt Design
In the w/o VT setting, the model directly generates both the reasoning path and the final answer based solely on the given query. In contrast, for the four visual thought paradigms (N-LANG, S-LANG, E-IMG, and G-IMG), a two-stage approach is employed. During the first stage, the model is prompted to generate a corresponding visual thought. During the second stage, this visual thought, integrated with the original query, serves as additional context to steer the model toward deriving the rationale and the final answer.
@@ C.1.2 Prompt Design for N-LANG
At this stage, the model is tasked with extracting meaningful information and transforming it into natural language, forming a foundation for visual thought and enabling downstream reasoning. The specific prompts employed in our experiments are listed below:
@@ C.1.4 Prompt Design for E-IMG
At this stage, E-IMG employs vision tools (e.g., Grounding-DINO [24], Semantic-SAM [20], DepthAnything [48]) to edit images based on the action series. These processed images serve as extended inputs, facilitating the formation of visual thoughts in the subsequent reasoning stage. The prompts employed in our experiments are listed below:
@@ C.1.5 Prompt Design for G-IMG
At this stage, G-IMG employs the vision model DALL·E 3 [1] to generate images based on tailored image generation prompts. These images embody the model’s visual thoughts and are used as additional inputs in the subsequent reasoning phase. The prompts employed in our experiments are listed below:
@@
C.2 Core Factors Evaluation Mechanism
@@
  • Image Relevance: This criterion assesses the extent to which the visual thoughts aligns with the semantic content and context of the input image. A highly relevant visual thoughts output should accurately reflect the objects, relationships, or scenes depicted in the image, ensuring that the generated content stays consistent with the source material.
@@ D.1 Visual Attention Analysis
In this section, we utilize LLaVA-1.5-7B [21] as the experiment model to analyze the attention distribution during its reasoning process and randomly select one sample from the ScienceQA [25] as a case study. Specifically, we compare two input settings: (1) a multimodal query alone, and (2) a multimodal query with visual thought. For each setting, we obtain the attention weights from the last token to the input image and the visual thought in each transformer layer.
@@
where $\bar{a}^{(l)}_{\text{last}\rightarrow k}$ represented the target extracted attention weight, $last$ indicates the index of the last token, $k$ indicates the index of the target token, and $H$ represents the number of the head.
As illustrated in Figure 7, the results indicate that the inclusion of visual thoughts leads to a shift in attention from the original image toward the visual thought. Moreover, the visual thought continues to carry and transmit visual information into deeper layers of the transformer.
@@ D.2.1 Attention Blocking Analysis
We manually block the information flow between specific tokens by setting the attention mask between them in selected transformer layers to –inf [40, 50]. As shown in Figure 8, blocking the information flow between the query and the visual thought leads the model to choose an incorrect option that it would otherwise have predicted correctly. This result emphasizes the crucial influence of information flow between the query and the visual thought on final answer prediction.
@@ D.2.2 Saliency-Based Information Flow Analysis
We compute saliency scores [36] to evaluate the relative importance of different information flows for answer prediction [40, 50]. The computing process utilize Taylor expansion [29] for attention matrix elements:
@@
Here, $\mathcal{L}(x)$ represents the loss function, and $x$ denotes the input; $A_{h,l}$ represents the value for the $h$-th attention head of the $l$-th layer in attention matrix; while $I_{l}$ denotes the saliency matrix, produced by averaging the values of all attention heads in the $l$-th layer, representing the significance of information flow.
@@
Appendix E Ethical Considerations

### v2

@@ Appendix A Experiment Setup
We conduct all our experiments using four models, including LLaVA-1.5 [21], Qwen2-VL [42], GPT-4o-mini [32], and GPT-4o [32]. For the GPT series models, we adjust the temperature parameter within [0,2]; for the open-source models, we adjust the temperature parameter within the range [0,2]. In addition, all open source models complete inference on 2 A6000 48G.
@@ B.1 Visual Thought Setting
Figure 11: Results of Direct and w/o Visual Thought prompting on GPT-4o.
@@ B.2 Image Classification
In this section we investigate how the complexity of an image’s textual description influences the effectiveness of visual reasoning. We first prompt the model to produce captions that are as precise as possible, and we quantify description difficulty by the number of tokens in each caption. Consistent with commonsense expectations and our manual annotations, images depicting spot-the-difference scenarios are the most challenging to describe and are therefore labeled Hard‑to‑Describe. Tangram‑related images present a moderate level of difficulty and are labeled Medium‑to‑Describe. By contrast, images showing mathematical functions detected through optical character recognition (OCR) require the fewest tokens and are labeled Easy‑to‑Describe.
B.3 Additional Discussion on w/o Visual Thoughts
To better substantiate that the performance degradation observed in the w/o Visual Thought scheme proposed in Figure  4 (a) does not merely stem from contextual perturbations, we compare direct prompting with GPT-4o to generate responses (Direct) against explicitly instructing it to avoid producing any visual content (w/o VT prompting). As shown in Figure  11, the results obtained by w/o VT prompting demonstrate a noticeable performance drop when visual information is entirely omitted by the model, even though the surrounding context remains complete.
B.4 More Results
Beyond the exploration in Section 3, we further investigate the effectiveness of incorporating visual thought into pure-text problems that require spatial imagination. In this setting, visual thought serves to convey visual information from the imagined spatial domain that is necessary for solving the textual problems. Specifically, we select tasks including chess, function, and graph from IsoBench and use Qwen2-VL as the experimental model. We employ VisualSketchpad[15] to generate reasoning processes that include image-form visual thought, keeping all other settings the same.
@@ Appendix C The Categories of Visual Thoughts
In this part of the experiment, we evaluate the performance of the four types of visual thoughts using four models, including LLaVA-1.5 [21], Qwen2-VL [42], GPT-4o-mini [32], and GPT-4o [32].
@@ C.1 Prompt Design
In the w/o VT setting, the model directly generates both the reasoning path and the final answer based solely on the given query. In contrast, for the four visual thought paradigms (N-LANG, S-LANG, E-IMG, and G-IMG), a two-stage framework is employed. In the first stage, the model is prompted to generate a corresponding visual thought. In the second stage, this visual thought, combined with the original query, serves as additional context to guide the model in deriving the reasoning path and the final answer.
@@ C.1.2 Prompt Design for N-LANG
At this stage, the model is tasked with extracting meaningful information and transforming it into natural language, forming a foundation for visual thought and enabling downstream reasoning. The specific prompts used in our experiments are as follows:
@@ C.1.4 Prompt Design for E-IMG
At this stage, E-IMG employs vision tools (e.g., Grounding-DINO [24], Semantic-SAM [20], DepthAnything [48]) to edit images based on the action series. These processed images serve as extended inputs, facilitating the formation of visual thoughts in the subsequent reasoning stage. The prompts used in our experiments are as follows:
@@ C.1.5 Prompt Design for G-IMG
At this stage, G-IMG employs the vision model DALL·E 3 [1] to generate images based on tailored image generation prompts. These images embody the model’s visual thoughts and are used as additional inputs in the subsequent reasoning phase. The prompts used in our experiments are as follows:
@@
Model MMVP V*Bench M3CoT CoMT AVG. position attributes physical social temporal deletion selection update w/o VT 74.33 53.95 54.78 88.89 76.86 79.67 26.50 19.50 37.00 56.83 N-LANG 85.33 57.89 63.48 88.89 78.93 83.74 33.50 25.50 37.50 61.64 S-LANG 84.33 63.16 64.35 90.00 78.51 82.93 29.50 18.00 42.00 61.42 E-IMG 83.00 59.21 65.22 90.00 78.10 86.18 34.00 28.50 50.00 63.80 G-IMG 78.00 59.21 59.13 92.22 78.93 86.18 33.50 28.50 46.50 62.46 w/o CoT 77.67 57.89 63.48 87.78 76.03 58.54 32.50 21.50 41.50 57.43
Table 3: Results of w/o CoT on GPT-4o.
C.2 More Results on w/o CoT
To gain a deeper understanding of the role of visual thoughts, we supplement the results in Table 1 with an additional vanilla baseline which is without any CoT reasoning, implemented using GPT-4o. As shown in the Table 3, we observe that the trends of removing Visual Thoughts (VT) and CoT are not consistent. For benchmarks that require more reasoning steps, such as M3CoT, which involves extensive multi-step reasoning, the performance of the w/o CoT variant significantly degrades.
C.3 Core Factors Evaluation Mechanism
@@
  • Image Relevance: This metric assesses the degree to which the visual thoughts aligns with the semantic content and context of the input image. A highly relevant visual thoughts output should accurately reflect the objects, relationships, or scenes depicted in the image, ensuring that the generated content remains faithful to the source material.
@@ D.1 Visual Attention Analysis
In this section, we employ LLaVA-1.5-7B [21] as the experiment model to analyze the attention distribution during its reasoning process and randomly select one sample from the ScienceQA [25] as a case study. Specifically, we compare two input settings: (1) a multimodal query alone, and (2) a multimodal query with visual thought. For each setting, we extract the attention weights from the last token to the input image and the visual thought in each transformer layer.
@@
where $\bar{a}^{(l)}_{\text{last}\rightarrow k}$ represented the target extracted attention weight, lastlast represents the index of the last token, $k$ represents the index of the target token, and $H$ represents the number of the head.
As illustrated in Figure 8, we observe that the inclusion of visual thoughts leads to a shift in attention from the original image toward the visual thought. Moreover, the visual thought continues to carry and transmit visual information into deeper layers of the transformer.
@@ D.2.1 Attention Blocking Analysis
We manually block the information flow between specific tokens by setting the attention mask between them in selected transformer layers to –inf [40, 50]. As shown in Figure 9, blocking the information flow between the query and the visual thought leads the model to choose an incorrect option that it would otherwise have predicted correctly. This result highlights the critical influence of information flow between the query and the visual thought on final answer prediction.
@@ D.2.2 Saliency-Based Information Flow Analysis
We compute saliency scores [36] to evaluate the relative importance of different information flows for answer prediction [40, 50]. The computing process utilize Taylor expansion [29] for each element of the attention matrix:
@@
Here, $\mathcal{L}(x)$ represents the loss function, and $x$ denotes the input; $A_{h,l}$ is the attention matrix value for the $h$-th attention head in the $l$-th layer. The saliency matrix $I_{l}$ for the $l$-th layer is obtained by averaging across all heads. The significance of information flow from the $j$-th token to the $i$-th token can be represented by $I_{l}(i,j)$.
@@
Model MMVP V*Bench M3CoT CoMT AVG. position attributes physical social temporal deletion selection update w/o VT 74.33 53.95 54.78 88.89 76.86 79.67 26.50 19.50 37.00 56.83 N-LANG(Maj@4) 85.00 57.89 63.48 88.89 78.93 83.74 33.50 25.50 37.50 61.60 S-LANG(Maj@4) 83.67 63.16 64.35 90.00 78.51 82.93 29.50 18.00 42.00 61.35 E-IMG(Maj@4) 82.33 59.21 65.22 92.22 78.10 86.18 33.50 28.00 49.00 63.75 G-IMG(Maj@4) 78.00 59.21 63.48 92.22 78.93 86.18 33.50 28.50 46.50 62.95 Diverse-VT 85.00 63.16 65.22 92.22 78.93 86.18 34.00 28.50 50.00 64.80
Table 4: Results of Diverse-VT on GPT-4o.
Appendix E Future Work
Based on the empirical analysis of visual thoughts, exploring ways to further enhance the performance of MCoT is a promising direction for future research. Here, inspired by self-consistency, we design a novel visual-thought-guided strategy that ensembles diverse visual thoughts reasoning paths, which are refered as diverse-VT. We implement diverse-VT on GPT-4o. As shown in Table 4, we observe that Diverse-VT achieves the best performance, demonstrating the feasibility of integrating multiple visual thoughts to enhance visual information transmission in MCoT, thereby improving reasoning performance.
Appendix F Ethical Considerations